# Less is More: Data Curation Matters in Scaling Speech Enhancement

### v1

@@ I Introduction
Speech enhancement (SE) [1] remains a cornerstone of robust speech processing systems, enabling critical applications from hearing aids to voice communication devices. Its core objective is to improve speech quality and intelligibility by removing undesired acoustic components from degraded signals. Thanks to the power of deep learning, neural network (NN)-based methods, including discriminative [2, 3, 4, 5, 6] and generative [7, 8, 9, 10] approaches, have made significant advances in speech enhancement over the past decade.
@@
Based on the findings of the prior works [12, 13, 15], this paper revisits the role of data scaling in speech enhancement from a new perspective. Acknowledging the known presence of low-quality samples within the “clean” labels of the URGENT2025 dataset (a key aspect of the challenge focused on noisy data utilization) [19], we first perform a preliminary analysis on its 2,500-hour training set. Given this context, our study focuses on investigating the trade-off between data scale and quality. We assume that these low-quality “clean” speech samples would have affected the scaling of the training data. To verify this hypothesis, we apply multiple non-intrusive evaluation metrics to the 2500-hour speech source of the URGENT2025 training set, and select the top 100-hour, 350-hour, and 700-hour subsets. We reviewed the effect of training data size on the performance of speech enhancement and compared it with the uniform random subset selection. Experiments in both discriminative and generative models show that curated subsets consistently outperform uniformly random samples of the same size. Crucially, we have made a surprising discovery that the system trained on a 700-hour optimal subset outperforms the one trained on the full 2,500-hour set in multiple evaluation metrics. This paper outlines our methodology and findings and proposes a data-centric data curation approach to advance future research on speech enhancement.
@@ II-C Data Curation Strategy
Metric EARS[24] Common Voice(ZH)[17] Default
@@ III Experiments
Model Training source non-intrusive metrics intrusive metrics
@@ III-B2 Generative Models
Figure 3: Impact of training data selection and scale on SE performance. Both discriminative BSRNN and generative BSRNN-Flow are evaluated. The X-axis is the amount of training speech source. The blue lines mark the systems trained with selected top-X-hour speech source; the orange lines mark the systems trained with uniformly sampled X-hour speech source. Intrusive and non-intrusive metrics are represented in different background colors.

### v2

@@ I Introduction
Speech enhancement (SE) [1] remains a cornerstone of robust speech processing systems, enabling critical applications from hearing aids to voice communication devices. Its core objective is to improve speech quality and intelligibility by removing undesired acoustic components from degraded signals. Thanks to the power of deep learning, neural network (NN)-based methods, including discriminative [2, 3, 4, 5, 6, 7] and generative [8, 9, 10, 11] approaches, have made significant advances in speech enhancement over the past decade.
@@
Based on the findings of the prior works [13, 14, 16], this paper revisits the role of data scaling in speech enhancement from a new perspective. Acknowledging the known presence of low-quality samples within the “clean” labels of the URGENT2025 dataset (a key aspect of the challenge focused on noisy data utilization) [20], we first perform a preliminary analysis on its 2,500-hour training set. Given this context, our study focuses on investigating the trade-off between data scale and quality. We assume that these low-quality “clean” speech samples would have affected the scaling of the training data. To verify this hypothesis, we apply multiple non-intrusive evaluation metrics to the 2500-hour speech source of the URGENT2025 training set, and select the top 100-hour, 350-hour, and 700-hour subsets. We reviewed the effect of training data size on the performance of speech enhancement and compared it with the uniform random subset selection. Experiments in both discriminative and generative models show that curated subsets consistently outperform uniformly random samples of the same size. Crucially, we discovered that the system trained on a 700-hour optimal subset outperforms the one trained on the full 2,500-hour set in multiple evaluation metrics. This paper outlines our methodology and findings and proposes a data-centric data curation approach to advance future research on speech enhancement.
@@ II-C Data Curation Strategy
Metric EARS[25] Common Voice(ZH)[18] Other
@@ III Experiments
Model
Training
source
non-intrusive metrics intrusive metrics
@@ III-B2 Generative Models
Figure 3: Impact of training data selection and scale on SE performance. Both discriminative BSRNN and generative BSRNN-Flow are evaluated. The X-axis is the amount of training speech source. The orange lines mark the systems trained with selected top-X-hour speech source; the blue lines mark the systems trained with uniformly sampled X-hour speech source. Intrusive and non-intrusive metrics are represented in different background colors.